OpenStudy (anonymous):

Please help me. 5^4*5^7 (1 point) A.)125^11 B.)15^11 C.)5^28 D.)5^11 (=

OpenStudy (anonymous):

D.)5^11

OpenStudy (anonymous):

are u positive?

OpenStudy (anonymous):

oh so what u did is left the 5 and added the exponents together right?

OpenStudy (nikato):

When ur multiplying things like those^, if u have the same base, u can just add the exponents

OpenStudy (anonymous):

a^4 times a^7 = a*a*a*a times a*a*a*a a*a*a= a*a*a*a a*a*a*a a*a*a

OpenStudy (anonymous):

Oh okay that makes since thanks. So what if the bases were different than each other?